OpenStudy (pottersheep):

Pleasee help. "A rug is needed to cover 2 thirds of the area of a room with dimensions 8 x 12. The uncovered part of the floor will form a border of uniform width around the rug. Determine the width of the border. Thanks so much !!!!!!!!!!

OpenStudy (anonymous):

 area of room=96 area of rug=64 64=(8-2x)(12-2x) 4x^2-40x+32=0 x^2-10x+8=0 x^2-10x+25=-8+25 (x-5)^2=17 x=5 + sqrt{17} this is the width of the border around the rug